OpenStudy (anonymous):

how to find the square root of 1089 the fast way? (without using calculator)

OpenStudy (anonymous):

its 33, theres no faster way than a calculator

OpenStudy (ash2326):

Write the factors of 1089 \[1089=9 \times 121\] \[1089=9 \times 11 \times 11\] \[1089=3 \times 3 \times 11 \times 11\] \[1089=3^2 \times 11^2\] so \[\sqrt {1089}=3 \times 11\]

OpenStudy (anonymous):

\[\sqrt{9}\sqrt{121}\] so thats 3 times 11 which is 33